AO Commissioning on 14 June 2011

Observer(s): G. Brusa, E. Pinna, A. Puglisi, S. Esposito, J. Hill, F. Quiros-Pacheo, J. Argomedo, K. Boutsia (LBTO)

Telescope Operator: S. Allanson (LBTO)

SW Support: N. Cushing (Tucson)

Telescope Support: J. Little (LBTO)

Instrument Support: T. Shih(LBTO)

Summary

We had a 'difficult' night characterized by two main issues: a) most of the acquire star instances failed and b) we had to re-fill PISCES dewar twice over the course of the night. Most of the night was lost troubleshooting the two mentioned issues. At the end of the night it was decided to replace the ASM Switch BCU with the spare in an attempt to fix the cause of multiple 'rips' from the ASM shell. The issue with PISCES was caused by an incorrect filling procedure, the correct procedure was identified and verified with Don McCarthy. Despite the issues we were able to complete two, relatively minor, tasks i.e.: verifying the WFS camera sensitivity and acquiring an ASM flat.

-- GuidoBrusa - 15 Jun 2011

Details

03:36 - we take test darks while they are changing the BCU unit
  • the darks have 0 values - this does not look right

03:47 iif_preset,Wref162, MODE=ACQUIRE, GS=0, ROTMODE=IDLE
  • to check routines and see if we still get 0 images
  • we can see the star but all the background is 0
  • we call doug miller - he sees the same thing on the images he is looking at after doing corquad.
  • we call craig and he sais that it seems that Pisces is warming up and this is why we get this background. We have two options - adjusting the background or refilling Pisces
  • Tim and Steve go to refill the camera

05:09 finish refilling camera

05:15 HBS failure - called J. Hill

05:24 iif_preset,Wref93new, MODE=ACQUIRE, GS=0, ROTMODE=IDLE

05:30 iif_preset,Wref93, MODE=ACE, GS=2, RS=0, PA=50
  • preset failed, because I used the old catalog which did not include the GS=2

05:33 iif_preset,Wref93new, MODE=ACE, GS=2, RS=0, PA=50
  • DIMM 0.9''

05:36 irc_runao
  • failed - the known bug that is opening the loop and the shell rips

05:42 irc_runao
  • failed - when we have such a failure we need to send again the preset

05:43 if_preset,Wref93new, MODE=ACE, GS=2, RS=0, PA=50

05:44 irc_runao
  • failed - the shell ripped

05:49 iif_preset,Wref93new, MODE=ACE, GS=2, RS=0, PA=50

05:50 irc_runao
  • failed - the shell ripped

05:55 iif_preset,Wref93new, MODE=ACE, GS=2, RS=0, PA=50

05:56 irc_runao - SUCCESSFUL!!!

06:27 iif_preset,Wref93new, MODE=ACE, GS=2, RS=0, PA=50

06:28 irc_runao - failed
  • loop opened without gioving any error at the iif
  • camera lense position error

06:55 iif_preset, FS23, MODE=ACQUIRE, GS=0, ROTMODE=IDLE - for standard
  • need pointing correction

06:59 iif_preset, FS23, MODE=ACQUIRE, GS=0, ROTMODE=IDLE

07:01 iif_preset, FS23, MODE=ACE, GS=15,RS=0, PA=50
  • preset failed because it cannot find the guide star

07:05 iif_preset, AS33, MODE=ACQUIRE, GS=0, ROTMODE=IDLE
  • we choose another star because the other one cannot find a guide star

07:11 No guide star found * going to a nearby BS star to collimate

07:20 iif_preset, AS33, MODE=ACQUIRE, GS=0, ROTMODE=IDLE

07:26 iif_preset, AS33, MODE=ACQUIRE, GS=5, ROTMODE=IDLE
  • changing guide star just to check what is wrong with the chart
  • still no guide star

07:28 iif_preset, FS23, MODE=ACQUIRE, GS=0, ROTMODE=IDLE
  • going back to the other star just to check if there is something wrong with the catalog or with something else...

07:33 iif_preset, FS23, MODE=ACE, GS=1, RS=0, PA=45
  • we found a guide star and we are taking images (pisces.20110614.0032.fits - *.0033.fits)
  • the detector has a wierd background and the temperature is going up
  • we are going to check pointing again just to verify that the pointing is correct
  • we took the probe out and obtaiben an image of the pointing star (pisces.20110614.0034.fits) - the star shows but the backgroung is low again,

07:50 the temperature is going up (80.4K) and we need to fill with LN again - Tim is going up

08:32 filling is over - temperature 75.14.

08:34 before moving to the next target the AO team does a test with the mirror to understand why the runao keep failing

08:52 iif_preset, AO749, MODE=ACQUIRE, GS=0, ROTMODE=IDLE

08:56 iif_preset, AO749, MODE=ACE, GS=14, RS=0, PA=150
  • we are trying a different guiding star because the target is too faint
  • guider seeing 1.1''
  • bad coordinates???

09:04 irc_runao - failed
  • thWaitMsg error(-5001): Timeout error - the shell ripped

09:22 re-start AGW and restart GCS (known bug??- feature??)
  • DIMM 1.0''

09:48 while the AO team is going on with their tests we take some pisces frames with different filtes starting at *.0038.fits
  • sky counts Ks 300cnts in 10sec / J 50-60cnts in 10 sec / J 150cnts in 100sec
  • .0053.fits - .0058.fits : closed loop
  • .0059.fits - .0062.fits : open loop

10:08 iif_preset, AO749, MODE=ACE, GS=14, RS=0, PA=150
  • reboot the AGW again

10:14 iif_preset, AO749, MODE=ACE, GS=7, RS=0, PA=150
  • the GCS was not restarted - T.O. cancelled the preset

10:16 iif_preset, AO749, MODE=ACE, GS=7, RS=0, PA=120
  • we stopped the collimation because the AO team wants to start their focus procedure from scratch

10:19 irc_runao
  • the loop closed but it did not arrive at the end of the procedure.
  • runao gave an error message
  • err = Resmsg[1] = Camera lens centering did not converge after 180 seconds

10:26 iif_preset, AO749, MODE=ACE, GS=7, RS=0, PA=120

10:27 irc_runao
  • failed again
  • err = Resmsg[0] = RunAO result status: Error
  • err = Resmsg[1] = err = Resmsg[2] = AcquireRef[57] failed

10:35 iif_preset, AO749, MODE=ACE, GS=7, RS=0, PA=120
  • guider seeing with the loop closed 0.5''
  • DIMM 0.8''

10:36 irc_runao
  • here they measured strehl

11:08 iif_preset, AO749, MODE=ACE, GS=7, RS=0, PA=120
  • stop guiding and wfs

11:10 irc_runao failes
  • err = Resmsg[0] = RunAO result status: Error
  • err = Resmsg[1] = err = Resmsg[2] = AcquireRef[69] failed
  • it stopped during the 10mode loop

11:16 iif_preset, AO751, MODE=ACQUIRE, GS=0, ROTMODE=IDLE

11:19 iif_preset, AO751, MODE=ACE, GS=16, RS=0, PA=140

11:22 irc_runao -failed
  • the algorithm did not work
  • err = Resmsg[1] = Timeout error on baysidey (-5001) TIMEOUT_ERROR

11:23 iif_preset, AO751, MODE=ACE, GS=16, RS=0, PA=140
  • preset failed becasue the AGW has to be restarted
  • they reboot the computer that controls the AGW

11:32 iif_preset, AO751, MODE=ACE, GS=16, RS=0, PA=140
  • the sky background is getting too high and they cannot really see the star

11:36 iif_preset, AS33, MODE=ACQUIRE, GS=0, ROTMODE=IDLE - for STD
  • check pointing
  • still no star for guiding!

11:45 iif_preset, BS9177, MODE=ACQUIRE, GS=0, ROTMODE=IDLE
  • we cannot guide because the background is too high
  • we take an image and it is completely out of focus
  • we give manual focus - we see the star and the J band has 2000 counts ( but backgrount 1500)
  • increase the exptime but we see no star - we use these frames as flats

12:40 darks with the lights on! - DO NOT USE
  • from 168 - with the lights off
  • it saturated at 40 sec...
  • during the 60 sec exposure... home the filter move it back to dark and it went to 5000 counts
  • probably we were looking through empty...
  • repeat darks tomorrow afternoon
  • during the last frame we turned on the light and half of the detector has 0 counts - one stripe at the bottom has normal counts (.0213.fits)
  • filter wheel had to be homed again - it was at the right filter
  • when a lot of filter changing is going on at some point the filter wheel fails...

sumary flats and standard

dwan flats filter
pisces.20110614.0080.hdr.fits - *.0085.fits J
pisces.20110614.0086.hdr.fits - *.0093.fits H
pisces.20110614.0100.hdr.fits - *.0105.fits Ks
pisces.20110614.0106.hdr.fits - *.0114.fits FeII 1.64 um
pisces.20110614.0116.hdr.fits - *.0126.fits Br-g 2.166 um
pisces.20110614.0129.hdr.fits - *.0137.fits H2 2.122 um\xA0

BS9177 but only ~500counts over background filter
pisces.20110614.0079.hdr.fits J
pisces.20110614.0094.hdr.fits - *.096.fits Ks

-- KonstantinaBoutsia - 14 Jun 2011

UT 6:03 testing the new version of auto-gain. This has a larger parameter window and should find the minimum of the slope signal rms better… (how is is exactly done?).

UT 6:06 optical loop data to check the magnitude of the star (the idea is to see if the WFS sensitivity is the same found during the last run). The magnitude is 7.5 vs. 7.7 las year.

UT 6:10 acquiring a new flat by averaging the AS position with CL. pisces.20110614.0030-31 shows saturated images of magH=4.4 with FeII

UT 8:34 we spent about 1/2 hour to refill PISCES dewar now the temperature is ~ 75 K.

UT 8:55 we are having troubles tracking on a 14.5 magnitude star, close to full moon.

UT 9:03 we found the star it was ~ 8" off, outside the FoV of the acquisition camera.

UT 9:10 the shell was 'ripped', it is not clear why.

UT 9:20 trying starting the loop manually but the shell 'ripped' again, it is not clear what happens.

UT 10:00 after some work on 'tweaking' the parameters of the automatic gain the shell 'ripped' again. Actuator #315 had a high current.

UT 10:52 we are able to close the loop but we now need to move the star in closed loop since we had to manually move the XY WFS stages in order to see the reference star.

UT 11:00 we found that some of the temperatures have values out of range, for instance the Water Inlet stays ~10C and then jumps in a few readings to up to 77 C...

UT 11:12 The shell 'ripped' again… when the loop was opened to move to another target.

-- GuidoBrusa - 15 Jun 2011
Topic revision: r7 - 19 Jun 2011, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback